OpenStudy (anonymous):

What is the side length of a cube with volume 4913 cm3?

OpenStudy (anonymous):

cube volume = L*L*L 4913 = L^3 can you do the rest?

OpenStudy (anonymous):

no

OpenStudy (♪chibiterasu):

I believe 17 . . .

OpenStudy (♪chibiterasu):

Yeah 17 :I

OpenStudy (anonymous):

so to do 4913 = L^3 take the cube root of both sides: \[\sqrt[3]{4913} = \sqrt[3]{L^3}\]